OpenStudy (lilah):

Sam sells necklaces for $10 each. Each necklace costs er $5 to make. She also has $5,000 in fixed costs per year for her jewelry business. How many necklaces must she sell in order to make a profit of $1,000 in one year? A. 1,200 B. 100 C. 600 D. 200

OpenStudy (lilah):

@jadeishere

OpenStudy (luffingsails):

Yearly Profit = SALES - COSTS Sales = 10 * x (where x is the number of necklaces sold in a year). Costs = 5000 (in fixed expenses) + 5 * x (where x is the same number of necklaces sold). Yearly Profit = 10x - (5000 + 5x) Set Yearly profit = 1000 and solve the equation.